# Beyond the Lens: Quantifying the Impact of Scientific Documentaries through Amazon Reviews

### v1

@@ Abstract.
Engaging the public with science is critical for a well-informed population. A popular method of scientific communication is documentaries. Once released, it can be difficult to assess the impact of such works on a large scale, due to the overhead required for in-depth audience feedback studies. In what follows, we overview our complementary approach to qualitative studies through quantitative impact and sentiment analysis of Amazon reviews for several scientific documentaries. In addition to developing a novel impact category taxonomy for this analysis, we release a dataset containing 1296 human-annotated sentences from 1043 Amazon reviews for six movies created in whole or part by a team of visualization designers who focus on cinematic presentations of scientific data. Using this data, we train and evaluate several machine learning and large language models, discussing their effectiveness and possible generalizability for documentaries beyond those focused on for this work. Themes are also extracted from our annotated dataset which, along with our large language model analysis, demonstrate a measure of the ability of scientific documentaries to engage with the public.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processingccs: Applied computing Media artsccs: Human-centered computing
@@ 1. Introduction
The impact of media was studied from the perspective of sociology, examining how documentaries reinforce or challenge cultural norms and societal structures (Rezapour and Diesner, 2017). The results show that documentaries can act as catalysts for social change by highlighting underrepresented issues and providing a platform for marginalized voices, thus fostering a more inclusive public dialogue around scientific and technological advancements (Bouzoubaa and Rezapour, 2024; Conrad-Pérez et al., 2022). These impacts extend beyond individual viewers to influence societal norms and values. By presenting scientific issues within relatable contexts, documentaries can alter public discourse, catalyze community actions, and even shift policy directions (Atakav, 2024). Moreover, the reach and accessibility of documentaries have been vastly expanded by advances in digital technology and online streaming platforms, which allow for unprecedented dissemination and engagement across diverse global audiences.
@@
Measuring the “impact” of scientific documentaries requires multimodal efforts. There are many decisions that designers of cinematic treatments must contend with in order to make an effective CSV presentation (Woodward et al., 2015; Borkiewicz et al., 2022; Jensen et al., 2023) and gauging audience attitudes and cognition can be difficult to do with precision, requiring one-on-one interviews and qualitative analysis (Chen, 2005; Cawthon and Moere, 2007; Buck, 2013; Fraser et al., 2012; Smith et al., 2015, 2017; Jensen et al., 2023, 2024). Thus, quantitative approaches can be used to complement in-depth qualitative analyses in order to better understand the impact of scientific media. With the emergence of digital platforms, viewers share their immediate reactions and detailed thoughts through online reviews, providing a rich dataset for analyzing public engagement and understanding (e.g., Appel et al., 2016). The interactivity provided by online platforms enhances the dissemination process, as viewers can discuss and share content, thereby amplifying the documentary’s reach and impact. This engagement is measurable through the analysis of online reviews and social media commentary, which serve as a feedback loop for content creators and scientists alike, indicating which aspects of their presentation meet the audience’s needs and which do not (Rezapour and Diesner, 2017; Bouzoubaa and Rezapour, 2024).
@@
Previous work showed that while online engagements offer insights into viewer reception, they often lack a systematic approach to quantify the depth and spectrum of influence exerted by documentaries on scientific understanding and behavioral change (Rezapour and Diesner, 2017). An additional shortcoming to prior work is that it often primarily relied on manual qualitative analysis, which is not only labor-intensive but also subject to bias and scalability issues (Bieniek-Tobasco et al., 2019; Atakav, 2024; Barrett and Leddy, 2008). We address this gap by computationally analyzing the impact of such scientific films using reviews and measuring the impact using a novel taxonomy of change and engagement in conjunction with a systematic approach to manual annotations. This approach employs natural language processing (NLP) techniques to systematically categorize and quantify viewer responses, thus providing a more scalable and objective measure of the documentaries’ effectiveness in enhancing scientific literacy and influencing viewers’ behavior. More specifically, we answer the following questions:
@@
To answer these questions, we extract reviews from Amazon for six prominent scientific documentaries. We apply text mining and sentiment analysis techniques to evaluate the tone, content, and thematic elements of viewer feedback. This data is further analyzed using our developed taxonomy, which categorizes impacts into “Shift in Cognition (C)”, “Attitudes Toward the Film (A)”, “Interest with Science Topic (S)”, “Impersonal report (I)”, and “Not applicable (N)”. Our data analysis shows that expressing different “Attitudes Toward the Film” is the most common Impact type reported in the reviews. The classification results suggest that large language model (LLM)-based classifiers, especially when the full context of the review is added to the prompt, are capable of successfully identifying Impact categories and their performance is generalizable to other datasets. Additionally, our thematic analysis of Impact groups demonstrates a diverse range of themes in the reviews. Viewers express positive, negative, or mixed feedback as a way of showing their attitude toward the film, while expressing environmental concerns or celebrating cosmic science are some of the themes related to shifts in cognition.
Our paper contributes to the field by (1) providing annotated data on impact, offering a dataset that has been carefully categorized and labeled to reflect the various types of influence that scientific documentaries have on public perception and interest; (2) introducing a novel taxonomy of engagement, which enables a structured analysis of how documentaries affect cognitive, affective, and interest dimensions of their audiences; and (3) employing advanced computational techniques, such as natural language processing and machine learning, to analyze viewer responses in a scalable and unbiased manner. These contributions not only enhance our understanding of the educational and societal impacts of scientific documentaries but also provide a framework for future research to build upon, improving the effectiveness and reach of science communication through visual media.
@@ 2. Related Work
Social media platforms are invaluable resources for understanding public sentiment and behavior, offering real-time access to users’ expressed opinions and emotions. By analyzing vast amounts of social media data, researchers can uncover trends, gauge audience reactions, and explore complex phenomena like consumer behavior and societal attitudes (El Alaoui et al., 2018; Cortis and Davis, 2021). Platforms like Twitter, Reddit, and YouTube provide large-scale datasets that enable studies to capture real-time feedback and public sentiment at an unprecedented scale in domains such as politics (Antypas et al., 2023; Yarchi et al., 2021), healthcare (Hussain and Al Ghawi, 2023; Arias et al., 2022), and education (Waheeb et al., 2022; Luo et al., 2020). For instance, sentiment analysis of tweets related to significant global events or media productions, such as the Netflix documentary Our Planet, reveals insights into public sentiment and media impact. This analysis helps creators and researchers understand audience engagement and the potential societal influence of content (Ravita Chahar, 2023). Similarly, Reddit’s discourse, known for its depth and variety, serves as a valuable source for exploring nuanced discussions and opinions (Liu and Zhang, 2012a; Deng et al., 2023; Guerra and Karakuş, 2023).
@@
Sentiment analysis has been central to leveraging social media data for understanding public opinion (Reyes-Menendez et al., 2018; Qian et al., 2022). Advances in NLP and machine learning have significantly improved sentiment classification tasks. Traditional approaches like Support Vector Machines (SVM) and K-Nearest Neighbors (KNN) have given way to deep learning methods, which excel in handling large, complex datasets (Bansal et al., 2022). Models such as Long Short-Term Memory (LSTM), Bidirectional LSTM (BiLSTM), and transformers like BERT and RoBERTa have achieved state-of-the-art performance in sentiment analysis (Kian Long Tan, 2022; Yasir Ali Solangi, 2019; Kapur and Harikrishnan, 2022). Recent comparisons of machine learning models for social media sentiment analysis reveal that deep learning approaches consistently outperform traditional algorithms. RoBERTa-LSTM hybrids, for instance, demonstrate exceptional accuracy and robustness in Twitter sentiment classification (Arif Ridho Lubis, 2023). These models’ ability to understand context and handle large datasets has made them indispensable for analyzing public opinion across diverse media platforms (Lyberius Ennio F. Taruc, 2024; Ravita Chahar, 2023; Niimi, 2024).
@@
With the emergence of Large Language Models (LLMs) such as GPT and BERT-based architectures, sentiment analysis has evolved further, leveraging pre-trained models to understand complex linguistic patterns and contextual subtleties. These models enable transfer learning, allowing fine-tuning on specific domains to achieve superior performance compared to traditional and earlier deep learning methods (Deng et al., 2023; Niimi, 2024). Studies emphasize the importance of domain-specific tuning of NLP models. For example, the application of sentiment analysis to social media reviews of environmental documentaries or films not only provides feedback on content reception but also sheds light on broader societal attitudes towards environmental issues (Acerbi et al., 2023).
Opinion mining goes a step further than sentiment analysis by focusing on specific aspects of text that inform user perspectives. This fine-grained analysis has been applied to social media and online discourse, uncovering detailed insights into user opinions on products, services, or cultural phenomena (Liu and Zhang, 2012a; Kapur and Harikrishnan, 2022; Gerard et al., 2023). For instance, (Kapur and Harikrishnan, 2022) employed BiLSTM combined with a random forest classifier to identify sentiments across multiple platforms, achieving high accuracy.
@@
Film reviews provide a specifically rich ground for sentiment analysis and opinion mining. They are rich in emotion, context, and critique, making them ideal for studying user opinions and their motivations. Existing research has predominantly focused on sentiment polarity, identifying whether reviews are positive, negative, or (occasionally) neutral. While effective, this binary/ternary classification often misses the granular details of user opinions (Mrabti et al., 2024; Palak Baid, 2017). Researchers have called for extending these efforts to include opinion mining, which examines the thematic and contextual dimensions of reviews. For instance, (Rezapour and Diesner, 2017) introduced “Impact” categories alongside traditional sentiment categories to analyze how specific aspects of films influence their public reception. This nuanced approach revealed critical factors driving engagement, such as emotional resonance and storytelling quality. In addition, (Yue et al., 2019) emphasizes the importance of incorporating fine-grained sentiment analysis to move beyond simple polarity classification. This involves analyzing not only whether reviews are positive or negative but also understanding the rich contextual and emotional dimensions that shape user opinions. By capturing these subtleties, researchers can uncover deeper insights into audience preferences and motivations, ultimately enabling more tailored and effective content strategies.
@@
Building on this body of work, our research aims to advance the application of NLP and machine learning techniques to film reviews, emphasizing both sentiment and impact analysis. By adopting a fine-grained approach, we seek to identify not only the overall sentiment but also the specific aspects of scientific films that shape public emotions and opinions. This holistic understanding will contribute to a deeper comprehension of user feedback, guiding content creators and industry stakeholders in optimizing their offerings.
@@ 3.1. Data Collection
A dataset of 1296 sentences was collected from a total of 1043 Amazon reviews of the six movies produced by a team of visualization designers who focus on cinematic presentations of scientific data (CSVT). Movies include: SuperTornado: Anatomy of a MegaDisaster (MagellanTV, 2015b), Birth of Planet Earth (MagellanTV, 2019), Solar Superstorms (TVF International, 2013; MagellanTV, 2015a), Seeing the Beginning of Time (MagellanTV, 2017), Space Junk (RLJ Entertainment, 2012), and The Jupiter Enigma (MagellanTV, 2018). In addition to the 1296 collected for annotation and model training, 400 sentences were collected from Amazon reviews of the the Hubble documentary (Toni Myers, 2010) for final model evaluation. Movies were chosen in collaboration with the [anonymous collaborators] of the CSVT as this team of visual designers expressed interest in quantifying the impact of their works.
Full reviews are downloaded using the selectorlib software package [^1] https://pypi.org/project/selectorlib/ and sentences are generated from full comments using spaCY’s (Honnibal and Montani, 2017) sentence tokenizer [^2] https://spacy.io/api/sentencizer. In addition to the full text, the date and rating of the review is also collected.
@@ 3.2.1. Taxonomy of Scientific Documentary Impact
In seeking evidence for constructivist learning (e.g., Bada and Olusegun, 2015), we aim to track written language that indicates a statement about cognition or a statement about engagement. As such, our impact categories are modeled off those used in the analysis of issue-focused documentaries (Rezapour and Diesner, 2017) with several updates for our set of scientific documentaries, namely any change or affirmation categories for behavior cognition or emotion (in Table 1, Rezapour and Diesner, 2017) have been replaced with the categories of “Shift in Cognition”, “Attitudes Toward the Film”, and “Interest with Science Topic”. More specifically, while internet comments are collected in too uncontrolled an environment to make any definitive conclusions about learning (e.g., Jensen, 2017), we can identify trends among contributors in their comments associated with learning. Our impact category “Shift in Cognition” represents audiences’ relative ability to identify terms, features, and concepts, up to and including metacognitive discussion of their own cognitive processes.
@@
3.3. Annotation Procedure
After reviewing the comments, we noticed that individual sentences within a comment can address different impacts and may express opposite sentiments. As a result, we opted to perform annotations at the sentence level. We used the Zooniverse [^3] https://www.zooniverse.org/ citizen science platform, one of the largest platforms for non-professionals to participate in the scientific process, for the annotation process.
Using Zooniverse’s default interface for textual data, we display a single sentence from a comment followed by the sentence within the full comment for context. This display and the first stage of the annotation process is depicted in Figure 1 in which the user is asked to select their first choice for the sentiment category (Positive, Negative, or Neutral). Here, ”first choice” is defined as the strongest sentiment expressed in the statement. If there are two equally strong sentiments expressed, then the sentiment that appears first is selected as the ”first choice” (the analysis of ”second-choice” selections are relegated to a future paper). Once the user has made their selection, they can select ”Done” to move to the next stage of the annotation (green button in Figure 1) or ”Done & Talk” to ask questions in the main forum (blue button in Figure 1). The annotator is then prompted to repeat the process with their first choice for impact category (categories listed in the first column of Table 1).
Annotation of each of the 1296 sentences occur in batches of $\approx$200 sentences, with each sentence in a batch “retired” once three users have completed annotations on the sentence. The breakdown of the number of annotated sentences per movie is shown in Table 2.
Figure 1. First stage of annotation process in which users are instructed to select their first choice for the sentiment category for each sentence. Each sentence (top) is shown with the context of the full review (bottom). Annotator moves to next stage of annotation with the green “Done” button (or blue “Done & Talk” button to post comments in the project forum). More information (comment URL and film title) are accessed with the circled “i” Info button, and annotation can be saved in the user’s personal collection for later reference with the heart Favorite button.
3.4. Data Cleaning and Analysis
After completing the annotation on Sentiment and Impact, we used Cohen’s Kappa (Cohen, 1960) to measure the degree of reliability and agreement between the annotators. For sentiment categories, the kappa score is relatively consistent across each category with 0.66 for “Positive” sentiments, 0.67 for “Neutral” sentiments, and 0.68 for “Negative” sentiments. For impact categories, the lowest Kappa measurement is for “Impersonal Report” (0.61) with higher measurements for “Interest in Science Topic”, “Shift in Cognition” and “Not applicable” (0.67) along with “Attitudes Toward the Film” (0.68).
Final sentiment and impact categories were decided by a majority vote. In $\approx$10% of the sentences, where no majority consensus was possible, all annotators discussed each sentence in a group and either a majority-consensus was reached, or the sentence was marked as ”no agreement”. This resulted in eight ”no agreement” sentences, which we exclude from our analysis. Finally, we excluded sentences that were only emojis or consisted of solely punctuation marks (e.g., ”:D”). The composition of the remaining 1286 sentences in sentiment and impact categories is shown as a percentages heatmap in Figure 2 with the per-media breakdown in raw numbers of sentiments of sentiment and impact category per film shown in Table 2.
@@ 3.4. Data Cleaning and Analysis
Shift in Cognition (C) Person encountered a new idea or way of thinking No change in way of thinking Person expressed disagreement with scientific concepts presented in film
“very informative about the development of tornados and the research behind their formation” “We need to reign in the junk orbiting our planet before e loose something more important than the satellite Iridium 23, like a human life!” “Unfortunately,HUGE lack of actual facts.”
Attitudes Toward the Film (A) Person indicates positive attributes about film attributes Person simply acknowledges film attributes Person expresses dislike of film attributes
@@
Interest with Science Topic (S) Person expresses positive interest in science topic presented in film Person expresses a connection with the science topic Person expresses disinterest in science topic
“Space intrigues me greatly, and this was an amazing program. ” “I’ve followed supercomputing since inception (and currently run some of their software).” “used to intrigue me as a child, but knowing reality and lies are now to be believed over actual evidence leaves me wanting real answers.”
Impersonal report (I) Person describes events in the film with positive connotations Person describes events in the film Person describes events in the film with negative connotations
@@
Not Applicable (N) Positive comment, but no specific relation to film or topic Comment has no specific relation to film or topic Negative comment, but no specific relation to film or topic

### v2

@@ Abstract.
Engaging the public with science is critical for a well-informed population. A popular method of scientific communication is documentaries. Once released, it can be difficult to assess the impact of such works on a large scale, due to the overhead required for in-depth audience feedback studies. In what follows, we overview our complementary approach to qualitative studies through quantitative impact and sentiment analysis of Amazon reviews for several scientific documentaries. In addition to developing a novel impact category taxonomy for this analysis, we release a dataset containing 1296 human-annotated sentences from 1043 Amazon reviews for six movies created in whole or part by the Advanced Visualization Lab (AVL). This interdisciplinary team is housed at the National Center for Supercomputing Applications and consists of visualization designers who focus on cinematic presentations of scientific data. Using this data, we train and evaluate several machine learning and large language models, discussing their effectiveness and possible generalizability for documentaries beyond those focused on for this work. Themes are also extracted from our annotated dataset which, along with our large language model analysis, demonstrate a measure of the ability of scientific documentaries to engage with the public.
@@
journalyear: 2025copyright: ccconference: 17th ACM Web Science Conference; May 20–24, 2025; New Brunswick, NJ, USAbooktitle: 17th ACM Web Science Conference (Websci ’25), May 20–24, 2025, New Brunswick, NJ, USAdoi: 10.1145/3717867.3717908isbn: 979-8-4007-1483-2/2025/05ccs: Computing methodologies Natural language processingccs: Applied computing Media artsccs: Human-centered computing
@@ 1. Introduction
The impact of media was studied from the perspective of sociology, examining how documentaries reinforce or challenge cultural norms and societal structures (Rezapour and Diesner, 2017; Diesner and Rezapour, 2015). The results show that documentaries can act as catalysts for social change by highlighting underrepresented issues and providing a platform for marginalized voices. This fosters a more inclusive public dialogue around scientific and technological advancements (Bouzoubaa and Rezapour, 2024; Conrad-Pérez et al., 2022; Diesner et al., 2016). These impacts extend beyond individual viewers to influence societal norms and values. By presenting scientific issues within relatable contexts, documentaries can alter public discourse, catalyze community actions, and even shift policy directions (Atakav, 2024). Moreover, the reach and accessibility of documentaries have been vastly expanded by advances in digital technology and online streaming platforms, which allow for unprecedented dissemination and engagement across varied global audiences.
Measuring the “impact” of scientific documentaries requires a multifaceted approach that integrates both qualitative and quantitative methods. Prior studies have demonstrated that filmmakers and designers of CSV presentations must navigate numerous creative and technical decisions to ensure their effectiveness (Woodward et al., 2015; Borkiewicz et al., 2022; Jensen et al., 2023). Traditionally, assessing audience attitudes and cognition has relied on in-depth qualitative techniques such as interviews and thematic analysis, which, while insightful, can be labor-intensive and subject to bias (Chen, 2005; Cawthon and Moere, 2007; Buck, 2013; Fraser et al., 2012; Smith et al., 2015, 2017; Jensen et al., 2023, 2024). Additionally, the reliance on manual qualitative analysis in prior research has posed challenges related to efficiency and reproducibility (Bieniek-Tobasco et al., 2019; Atakav, 2024; Barrett and Leddy, 2008).
@@
To address these limitations, quantitative approaches can serve as valuable complements, enhancing precision and scalability in impact assessments. By integrating quantitative methods, researchers can overcome the constraints of qualitative-only studies, leading to a more comprehensive and rigorous evaluation of scientific media’s influence. With the emergence of digital platforms, viewers share their immediate reactions and detailed thoughts through online reviews, providing a rich dataset for analyzing public engagement and understanding (e.g., Appel et al., 2016). The interactivity provided by online platforms enhances the dissemination process, as viewers can discuss and share content, thereby amplifying the documentary’s reach and impact. This engagement is measurable through the analysis of online reviews and social media commentary, which can serve as a feedback loop for content creators and scientists alike, indicating which aspects of their presentation meet the audience’s needs and which do not (Rezapour and Diesner, 2017; Bouzoubaa and Rezapour, 2024).
@@
We extend previous work by computationally analyzing the impact of scientific films using user-generated reviews and measuring the impact using a novel taxonomy of change and engagement in conjunction with a systematic approach to manual annotations. This approach employs natural language processing (NLP) techniques to systematically categorize and quantify viewer responses, thus providing a more scalable and objective measure of the documentaries’ effectiveness in enhancing scientific literacy and influencing viewers’ behavior. More specifically, we answer the following questions:
@@
To answer these questions, we extract reviews from Amazon for six prominent scientific documentaries and evaluate the tone, content, and thematic elements of viewer feedback. This data is further analyzed using our developed taxonomy, which categorizes impacts into “Shift in Cognition (C)”, “Attitudes Toward the Film (A)”, “Interest with Science Topic (S)”, “Impersonal report (I)”, and “Not applicable (N)”. Our data analysis shows that expressing different “Attitudes Toward the Film” is the most common Impact type reported in the reviews. The classification results suggest that large language model (LLM)-based classifiers, especially when the full context of the review is added to the prompt, are capable of successfully identifying Impact categories and their performance is generalizable to other datasets. Additionally, our thematic analysis of Impact groups demonstrates a diverse range of themes in the reviews. Viewers used mixed sentiments to show their attitude toward the film, while expressing environmental concerns or celebrating cosmic science were some of the themes related to shifts in cognition.
Our paper makes the following contributions: (1) providing annotated data on impact, offering a dataset that has been carefully categorized and labeled to reflect the various types of influence that scientific documentaries have on public perception and interest [^1] https://huggingface.co/datasets/scidoc/websci2025; (2) introducing a novel taxonomy of engagement, which enables a structured analysis of how documentaries affect cognitive, affective, and interest dimensions of their audiences; and (3) employ advanced computational techniques to analyze viewer responses in a scalable and unbiased manner. These contributions enhance our understanding of the educational and societal impacts of scientific documentaries. The paper also provides a framework for future research to build upon, improving the effectiveness and reach of science communication through visual media.
@@ 2. Related Work
Social media platforms are invaluable resources for understanding public sentiment and behavior, offering real-time access to users’ expressed opinions and emotions. By analyzing vast amounts of social media data, researchers can uncover trends, gauge audience reactions, and explore complex phenomena like consumer behavior and societal attitudes (El Alaoui et al., 2018; Cortis and Davis, 2021). Platforms like Twitter, Reddit, and YouTube provide large-scale datasets that enable studies to capture real-time feedback and public sentiment at an unprecedented scale in domains such as politics (Antypas et al., 2023; Yarchi et al., 2021; Rezapour et al., 2017), healthcare (Hussain and Al Ghawi, 2023; Arias et al., 2022), and education (Waheeb et al., 2022; Luo et al., 2020). For instance, sentiment analysis of tweets related to significant global events or media productions, such as the Netflix documentary Our Planet, reveals insights into public sentiment and media impact. This analysis helps creators and researchers understand audience engagement and the potential societal influence of content (Ravita Chahar, 2023). Similarly, Reddit’s discourse, known for its depth and variety, serves as a valuable source for exploring nuanced discussions and opinions (Liu and Zhang, 2012a; Deng et al., 2023; Guerra and Karakuş, 2023; Bouzoubaa and Rezapour, 2024).
@@
Sentiment analysis has been central to leveraging social media data for understanding public opinion (Reyes-Menendez et al., 2018; Qian et al., 2022). Advances in NLP and machine learning have significantly improved sentiment classification tasks. Traditional approaches like Support Vector Machines (SVM) and K-Nearest Neighbors (KNN) have given way to deep learning methods, which excel in handling large, complex datasets (Bansal et al., 2022). Models such as Long Short-Term Memory (LSTM), Bidirectional LSTM (BiLSTM), and transformers like BERT and RoBERTa have achieved state-of-the-art performance in sentiment analysis (Kian Long Tan, 2022; Yasir Ali Solangi, 2019; Kapur and Harikrishnan, 2022). Recent comparisons of machine learning models for social media sentiment analysis reveal that deep learning approaches consistently outperform traditional algorithms. RoBERTa-LSTM hybrids, for instance, demonstrate exceptional accuracy and robustness in Twitter sentiment classification (Arif Ridho Lubis, 2023). These models’ ability to understand the context and handle large datasets has made them indispensable for analyzing public opinion across diverse media platforms (Lyberius Ennio F. Taruc, 2024; Ravita Chahar, 2023; Niimi, 2024).
@@
With the emergence of LLMs such as GPT and BERT-based architectures, sentiment analysis has evolved further, leveraging pre-trained models to understand complex linguistic patterns and contextual subtleties(Sayeed et al., 2023; Chang et al., 2024). These models enable transfer learning, allowing fine-tuning on specific domains to achieve superior performance compared to traditional and earlier deep learning methods (Deng et al., 2023; Niimi, 2024). Studies emphasize the importance of domain-specific tuning of NLP models. For example, the application of sentiment analysis to social media reviews of environmental documentaries or films not only provides feedback on content reception but also sheds light on broader societal attitudes towards environmental issues (Acerbi et al., 2023). Opinion mining goes a step further than sentiment analysis by focusing on specific aspects of text that inform user perspectives. This fine-grained analysis has been applied to social media and online discourse, uncovering detailed insights into user opinions on products, services, or cultural phenomena (Liu and Zhang, 2012a; Kapur and Harikrishnan, 2022; Gerard et al., 2023). For instance, (Kapur and Harikrishnan, 2022) employed BiLSTM combined with a random forest classifier to identify sentiments across multiple platforms, achieving high accuracy.
@@
Film reviews provide a specifically rich ground for sentiment analysis and opinion mining (Malini and Sunitha, 2019). They are rich in emotion, context, and critique, making them ideal for studying user opinions and their motivations. Existing research has predominantly focused on sentiment polarity, identifying whether reviews are positive, negative, or neutral. While effective, this binary/ternary classification often misses the granular details of user opinions (Mrabti et al., 2024; Palak Baid, 2017). Researchers have called for extending these efforts to include opinion mining, which examines the thematic and contextual dimensions of reviews. For instance, (Rezapour and Diesner, 2017) introduced “Impact” categories to analyze how specific aspects of films influence people’s cognition and attitude. This nuanced approach revealed critical factors driving engagement, such as emotional resonance and storytelling quality. In addition, (Yue et al., 2019) emphasizes the importance of incorporating fine-grained sentiment analysis to move beyond simple polarity classification. This involves analyzing not only whether reviews are positive or negative but also understanding the rich contextual and emotional dimensions that shape user opinions. By capturing these subtleties, researchers can uncover deeper insights into audience preferences and motivations, ultimately enabling more tailored and effective content strategies.
@@
Building on this body of work, by adopting a fine-grained approach, we seek to identify not only the overall sentiment but also the specific aspects of scientific films that shape public emotions and opinions. This holistic understanding will contribute to a deeper comprehension of user feedback, guiding content creators and industry stakeholders in optimizing their offerings.
@@ 3.1. Data Collection
A dataset of 1296 sentences was collected from a total of 1043 Amazon reviews of the six movies produced by a team of visualization designers who focus on cinematic presentations of scientific data (CSVT). The mean word count per review is 11.2 with a standard deviation of 9.0 words, and the mean word length across all reviews is 4.6 characters with a standard deviation of 2.7 characters. Movies include: SuperTornado: Anatomy of a MegaDisaster (MagellanTV, 2015b), Birth of Planet Earth (MagellanTV, 2019), Solar Superstorms (TVF International, 2013; MagellanTV, 2015a), Seeing the Beginning of Time (MagellanTV, 2017), Space Junk (RLJ Entertainment, 2012), and The Jupiter Enigma (MagellanTV, 2018). In addition to the 1296 collected for annotation and model training, 400 sentences were collected from Amazon reviews of the Hubble documentary (Toni Myers, 2010) for final model evaluation. Movies were selected in collaboration with the AVL as this team of visual designers expressed interest in quantifying the impact of their works. Full reviews were downloaded using the selectorlib software package [^2] https://pypi.org/project/selectorlib/. In addition to the full text, the date and rating of the review were also collected.
@@ 3.2.1. Taxonomy of Scientific Documentary Impact
In seeking evidence for constructivist learning (e.g., Bada and Olusegun, 2015), we aim to track written language that indicates a statement about cognition or a statement about engagement (Table 1). As such, our impact categories are modeled off those used in the analysis of issue-focused documentaries (Rezapour and Diesner, 2017) with several updates for our set of scientific documentaries(in Table 1, Rezapour and Diesner, 2017). More specifically, while internet comments are collected in too uncontrolled an environment to make any definitive conclusions about learning (e.g., Jensen, 2017), we can identify trends among contributors in their comments associated with learning. Our impact category “Shift in Cognition” represents audiences’ relative ability to identify terms, features, and concepts, up to and including metacognitive discussion of their own cognitive processes.
@@
Shift in Cognition (C): ability to identify terms, features, and concepts, up to and including metacognitive elements Person encountered a new idea or way of thinking No change in way of thinking Person expressed disagreement with scientific concepts presented in film
“very informative about the development of tornados and the research behind their formation” “We need to reign in the junk orbiting our planet before we lose something more important than the satellite Iridium 23, like a human life!” “Unfortunately, HUGE lack of actual facts.”
Attitudes Toward the Film (A): evaluations of film technical and artistic attributes Person indicates positive attributes about film attributes Person simply acknowledges film attributes Person expresses dislike of film attributes
@@
Interest with Science Topic (S): evaluations and connections to film scientific content Person expresses positive interest in science topic presented in film Person expresses a connection with the science topic Person expresses disinterest in science topic
“Space intrigues me greatly, and this was an amazing program.” “I’ve followed supercomputing since inception (and currently run some of their software).” “used to intrigue me as a child, but knowing reality now, I seek real answers over lies.”
Impersonal report (I): descriptions of events in a film without evaluation Person describes events in the film with positive connotations Person describes events in the film Person describes events in the film with negative connotations
@@
Not Applicable (N): statements with no specific relation to film or science topic Positive comment, but no specific relation to film or topic Comment has no specific relation to film or topic Negative comment, but no specific relation to film or topic
@@
3.3. Annotation Procedure
After reviewing a sample of reviews, we noticed that individual sentences within a review can address different impacts and may express opposite sentiments. As a result, we opted to perform annotations at the sentence level. We used the Zooniverse [^3] https://www.zooniverse.org/ citizen science platform, one of the largest platforms for non-professionals to participate in the scientific process, for the annotation process.
Using Zooniverse’s default interface for textual data, we displayed a single sentence from a review followed by the sentence within the full review for context. This display and the first stage of the annotation process is depicted in Figure 2 in the Appendix in which the user is asked to select their first choice for the sentiment category (Positive, Negative, or Neutral).
Annotation of each of the 1296 sentences occured in batches of $\approx$200 sentences, with each sentence in a batch “retired” once three users have completed annotations on the sentence. The breakdown of the number of annotated sentences per movie is shown in Table 2.
3.4. Data Cleaning and Analysis
After completing the annotation on Sentiment and Impact, we used Cohen’s Kappa (Cohen, 1960) to measure the degree of reliability and agreement between the annotators. For sentiment categories, the kappa score was relatively consistent across each category with 0.66 for “Positive” sentiments, 0.67 for “Neutral” sentiments, and 0.68 for “Negative” sentiments. For impact categories, the lowest Kappa measurement was for “Impersonal Report” (0.61) with higher measurements for “Interest in Science Topic”, “Shift in Cognition” and “Not applicable” (0.67) along with “Attitudes Toward the Film” (0.68).
Final sentiment and impact categories were decided by a majority vote. In $\approx$10of the sentences, where no majority consensus was possible, all annotators discussed each sentence in a group and either a majority consensus was reached, or the sentence was marked as “no agreement”. This resulted in eight “no agreement” sentences, which we excluded from our analysis. Finally, we excluded sentences that were only emojis or consisted of solely punctuation marks (e.g., “:D”). The composition of the remaining 1286 sentences in sentiment and impact categories is shown as a percentages heatmap in Figure 1 with the per-media breakdown in raw numbers of sentiments of sentiment and impact category per film shown in Table 2.